The perimeter of an equilateral triangle is 77 inches more than the perimeter of a square, and the side of the triangle is 55 inches longer than the side of the square. Find the side of the triangle.​ (Hint: An equilateral triangle has three sides the same​ length.
Answer by ikleyn(52787)

.
The perimeter of an equilateral triangle is 77 inches more than the perimeter of a square, and the side of the triangle is 55 inches longer than the side of the square. Find the side of the triangle.​ (Hint: An equilateral triangle has three sides the same​ length.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

From the condition, you have this system of equations:

3t = 4s + 77,     (1)
 t = s + 55,      (2)

where t = the length of the side of the triangle, s is that of square.

To solve it, simply substitute the expression t = s + 5 of (2) into equation (1). You will get an equation for one single unknown

3*(s+55) = 4s + 77,

which is easy to solve:

s = 165 - 77 = 88.

Then t = s + 55 = 88 + 55 = 143.